OpenStudy (anonymous):

Find the slope of the line when the points are (7.0) and (-3,7)

myininaya (myininaya):

\[\frac{ \Delta y}{\Delta x}=\frac{y_1-y_2}{x_1-x_2}\] you can do either \[\frac{7-0}{-3-7}\] or you can do \[\frac{0-7}{7-(-3)}\] to find the slope

OpenStudy (anonymous):

I am confused

myininaya (myininaya):

tell me exactly where you confusion occurs please

OpenStudy (anonymous):

how to solve for the slope

myininaya (myininaya):

ok you do the change in y divided by the change in x

jimthompson5910 (jim_thompson5910):

Slope \[m=\frac{y_2-y_1}{x_2-x_1}\] \[m=\frac{7-0}{-3-7}\] \[m=\frac{7}{-10}\] \[m=-\frac{7}{10}\]

myininaya (myininaya):

oh where you having problem with the subtraction on top and on bottom\

myininaya (myininaya):

were*